# Effect of dark matter on the shadow of a distorted and deformed compact object

### v1

@@
Abstract
@@
This work investigates observational properties, namely the shadow and photon ring structure, of emission profiles originated near compact objects. In particular, we consider a distorted and deformed compact object characterized by quadrupoles and surrounded by an optically thin and geometrically thin accretion disk with different emission profiles modeled by the Johnson’s Standard-Unbound (SU) distribution in the reference frame of the emitter. Under these assumptions, we produce the observed intensity profiles and shadow images for a face-on observer. Our results indicate that, due to the fact that modifications of the quadrupole parameters affect the radius of the innermost stable circular orbit (ISCO) and the photon sphere (PS), the observed shadow images and their properties are significantly influenced by the quadrupole parameters and emission profiles. Furthermore, we analyze the impact of the presence of a dark matter halo in the observational imprints considered and verify that both the increase in the matter contained in the halo or decrease in the length-scale of the halo lead to an increase in the size of the observed shadow. Our results indicate potential degeneracies between the observational features of distorded and deformed compact objects with those of spherically symmetric blackholes, which could be assessed by a comparison with the current and future generation of optical experiments in gravitational physics.
preprint: APS/123-QED
I Introduction
@@
The Event Horizon Telescope (EHT) Collaboration recently reported 1.3 mm Very Long Baseline Interferometry (VLBI) observations of the center of both the M87* galaxy and the object Sgr A* in the core of our own galaxy. The EHT Collaboration has reported the observation of an annular feature centered on M87* with a typical radius of $\sim 21\mu as$. The brightness along the ring is asymmetric mostly because of the Doppler boosting from the matter orbiting around the jet axis, which is inclined by 17 degrees relative to the line of sight [1, 2, 3, 4, 5, 6, 7, 8]. The ring diameter of the central source of our Milky Way, Sgr A*, located at a distance of $8kpc$ is reported as $\sim 50\mu as$, with an image that is dominated by a bright, thick ring with a diameter of $51.8\pm 2.3\mu as$. This ring has a modest azimuthal brightness asymmetry and a comparatively dim interior. Preferably the dominant model of the accretion disk that reproduces these observations is a magnetically arrested disk with inclination $i\leq 30$ degrees. [9, 10, 11, 12, 13].
@@
The shadow and surrounding photon ring involve a special curve known as the apparent boundary [14], and it is well known as the critical curve on the image plane. A light ray with an impact parameter equal to that of the critical curve asymptotically approaches a bound photon orbit when traced backwards from its observation by a distant observer. The existence of such a critical curve produces an observable photon ring: a bright and narrow ring-shaped feature near the critical curve due to the emission of radiation by the optically thin matter from the region the central object, whose photons travel close to the critical curve before reaching the observer [15, 16, 17].
@@
Additionally, the light rays associated with the photon ring typically orbit the central object multiple times before reaching the observer. This can lead to a significant increase in brightness that may continue to grow without limit in the absence of absorption, as the optical path lengths near the critical curve become increasingly long; nevertheless, the specific form of the visible photon ring is not influenced by the astronomical source profile [18]. Therefore it can be used as a probe to test general relativity and models for compact objects. In the absence of rotation, one obtains a single circular critical curve corresponding to an extremum of the effective potential, which is called a photon sphere [19].
@@
Separating the contributions of the background geometry and the influence of the surrounding flow in shadow images has become a critical challenge in the field. In the case of a spherical accretion configuration, the size of the observed shadow for a Schwarzschild black hole is closely influenced only by the spacetime geometry, without being affected by the emission profile [20]. However, considering an optically and geometrically thin disk around a compact object, later studies indicate that the black hole image is composed of a dark region and a bright region consisting of several components: direct emission, lensing rings, and photon rings, and the size of the shadow depends on the location of the inner edge of the accretion disk, and the intensity of the observed luminosity [21]. Those important features may be regarded as a way to distinguish between black holes and black hole mimickers in the universe [22, 23, 24] In fact, the brightness and width of the lensing ring, as well as the corresponding luminosity, depend on the background geometry and the emission features of the accretion disk surrounding the compact object, and the observed image is dominated by the direct component [25]. Nonetheless, this is only true for axial observations, as the direct and lensed contributions lose their circular structure for inclined observations [26, 27, 28].
@@
If the spacetime under consideration features an innermost stable circular orbit (ISCO), i.e., a radius below which circular orbits become unstable, this radius accurately predicts where the inner edge of the accretion disk lies [29]. However, in models of optically thin and geometrically thick accretions disks, for which the inner edge of the disk can penetrate inside the critical curve, then the minimum size of the shadow is not necessarily limited to the critical curve and can be strongly reduced [30, 31]. One of the main advantages of shadow observations is to provide a framework with which to probe deviations with respect to the Kerr or Schwarzschild solutions. However, the limited precision of the first set of observations does not allow for a precise identification of the underlying geometry [32, 33, 34]. Therefore, it is difficult to underestimate theoretical efforts to calculate forms and properties of shadows produced by black holes or other alternative compact objects in astrophysical environments. Besides these two well-studied Schwarzschild and Kerr hypotheses, at the theoretical level several alternative models for compact objects have been proposed, featuring a wide range of properties that could potentially fill the gap in the interpretation and understanding of the observational data [35].
@@
In this work, we are interested in studying the shadow properties in the background of a distorted and deformed compact object due to quadrupole moments. This solution is a generalization of the $\rm q$-metric previously analyzed in other publications e.g., [36, 37]. We show that moderate changes in the quadrupole moment could lead to significant alterations in the properties of the shadow. On one hand, we see that the radius of the photon ring, the innermost stable circular orbit (ISCO), and optical appearance can be different depending on the parameters of the geometry and intensity profile. On the other hand, we can deduce information about the source by its observed intensity profile. The latter has a crucial importance on the test of the nature of compact objects and, therefore, on the test of the strong field regime in the general theory of relativity.
@@
Another important issue in astrophysics is the observation of abnormal high velocities in the outer region of galaxies due to the surrounding invisible dark matter [38, 39]. The gravitational tidal forces induced by dark matter influence the propagation of light. Therefore, the presence of dark matter can be analyzed through the properties of the shadow via the assumption of different equations of state e.g. [40, 41, 42]. However, in these studies the results are model-dependent. To explore the effect of dark matter, we consider its behavior as an effective mass, which leads to modifications in the background geometry e.g. [43, 44, 45, 46].
@@
The paper is organized as follows: in Section II, we introduce the generalized $\rm q$-metric; in Section III we review the geodesic motion of timelike and null particles in this background, we introduce the accretion disk models, and we produce the observed shadow images and intensity profiles; in Section IV we introduce the dark matter model as a modification of the radial mass function and obtain the corresponding observed shadow images and intensity profiles; and in Section V we summarize our results and trace prospects for future work. Throughout the paper, we use a system of geometric units in which $G=1=c$ unless stated otherwise.
II Generalized $\rm q$-metric spacetime
@@
One of the generalizations of the Schwarzschild spacetime recurring to the Weyl’s family of solutions [47] is the $\rm q$-metric. This asymptotically flat solution of the Einstein field equations represents the exterior gravitational field of an isolated static and axisymmetric mass distribution containing a quadrupole moment, and it can be used to investigate the exterior fields of slightly deformed astrophysical objects in the strong-field regime [48]. This metric has been further generalized by relaxing the assumption of an isolated compact object, or equivalently asymptotic flatness, by considering the presence of an exterior distribution of mass in its vicinity [37]. The result of such a generalization resembles a magnetic surrounding [49] and it is also characterized by multipole moments. However, due to the subdominant effects of the higher multipole moments, it is appropriate to consider only the contributions to the metric up to the quadrupole moment for analyses in an astrophysical context. In the usual spherical coordinates $\left(t,r,\theta,\phi\right)$ the line element describing this generalized $\rm q$-metric is written as
@@
where $M$ is a parameter that plays the role of the mass of the central object, $\alpha$ and $\beta$ represent the multipole parameters, and the coordinates range in the intervals $t\in(-\infty,+\infty)$, $r\in(0,+\infty)$, $\theta\in[0,\pi]$, and $\phi\in[0,2\pi)$. Up to the quadrupole $\beta$, the external fields read as follows
@@
In the equatorial plane, described by $\theta=\pi/2$, the external fields above reduce to
@@
$\hat{\gamma}$ $=-2\beta(1+\alpha)\left(\frac{r}{M}-1\right)+\frac{\beta^{2}r}{4M}\left(\frac{r}{M}-2\right)\left(\left(\frac{r}{M}-1\right)^{2}-1\right)\,.$
Under the assumptions outlined above, this metric contains three free parameters: the deformation parameter $\alpha$ which is connected to the deformation of the compact object, the distortion $\beta$ which is connected to the external mass distribution, and the total mass $M$ that can be absorbed through a redefinition of the radial coordinate. In fact, the presence of a quadrupole can have a notable impact on the geometric properties of space-time when compared to the Schwarzschild solution see e.g. [50, 51]. Mathematically, the deformation parameter is bounded by $\alpha\in(-1,\infty)$, whereas the parameter $\beta$ in the equatorial plane depends on the value of $\alpha$ and it turned out to be significantly smaller than the latter. However, in an appropriate astrophysical context, other studies see e.g. [50, 51] have demonstrated that the most relevant range for $\alpha$ is $(-1,1)$; Additionally, in this work, we are interested in configurations that are approximately spherically symmetric, and consequently, quadrupoles should be relatively small in magnitude.This solution is the exact superposition of an oblate or prolate compact object, surrounded by an external distribution of matter. Recently, the impact of superposed fields in the motion of free test particles in this spacetime was studied [37] and it was shown how the positions of important astrophysical orbits, such as the innermost stable circular orbit (ISCO) for particles, and photon circular orbits, depend on the parameters of the metric. Moreover, from an astrophysical perspective, the negative values of quadrupoles are rather counter-intuitive since they describe prolated compact objects, while positive values correspond to ablated compact objects that are vastly observed in the universe.
III Shadows and light rings

### v2

@@
11institutetext: University of Bremen, Center of Applied Space Technology and Microgravity (ZARM), 28359 Germany 22institutetext: Waterloo Centre for Astrophysics, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada 33institutetext: Department of Physics and Astronomy, University of Waterloo, 200 University Avenue West, N2L 3G1 Waterloo, Canada 44institutetext: Perimeter Institute for Theoretical Physics, 31 Caroline Street North, Waterloo, Canada
44email: s3faraji@uwaterloo.ca
55institutetext: Institute of Theoretical Physics and Astrophysics, University of Gdańsk, Jana Bażyńskiego 8, 80-309 Gdańsk, Poland 66institutetext: Institute of Physics, University of Tartu, W. Ostwaldi 1, 50411 Tartu, Estonia
66email: joaoluis92@gmail.com
Shadow of quadrupole-deformed compact objects in a local dark matter shell
Shokoufe Faraji 11223344    João Luís Rosa 5566
(Received ……; accepted …..)
@@
This work investigates observational properties, namely the shadow and photon ring structure, of emission profiles originating near compact objects. In particular, we consider a distorted and deformed compact object characterised by two quadrupole parameters and surrounded by an optically thin and geometrically thin accretion disk with different emission profiles modelled by Johnson’s Standard-Unbound (SU) distribution in the reference frame of the emitter. Under these assumptions, we produce the observed intensity profiles and shadow images for a face-on observer. Our results indicate that, due to the fact that modifications of the quadrupole parameters affect the radius of the innermost stable circular orbit (ISCO) and the unstable photon orbits on the equatorial plane, the observed shadow images and their properties are significantly influenced by the quadrupole parameters and emission profiles. Furthermore, we analyse the impact of the presence of a dark matter halo in the observational imprints considered and verify that both the increase in the matter contained in the halo or the decrease in the length-scale of the halo lead to an increase in the size of the observed shadow. Our results indicate potential degeneracies between the observational features of distorted and deformed compact objects with those of spherically symmetric blackholes, which could be assessed by a comparison with the current and future generation of optical experiments in gravitational physics.
Key Words.:
Accretion, accretion disks – Black hole physics – Gravitation – Gravitational lensing: strong – Methods: numerical
1 Introduction
@@
The Event Horizon Telescope (EHT) Collaboration recently reported 1.3 mm Very Long Baseline Interferometry (VLBI) observations of the centre of both the M87* galaxy and the object Sgr A* in the core of our own galaxy. The EHT Collaboration has reported the observation of an annular feature centred on M87* with a typical radius of $\sim 21\mu as$. The brightness along the ring is asymmetric mostly because of the Doppler boosting from the matter orbiting around the jet axis, which is inclined by 17 degrees relative to the line of sight (Event Horizon Telescope Collaboration, 2019a, b, c, d, e, f, 2021a, 2021b). The ring diameter of the central source of our Milky Way, Sgr A*, located at a distance of $8kpc$ is reported as $\sim 50\mu as$, with an image that is dominated by a bright, thick ring with a diameter of $51.8\pm 2.3\mu as$. This ring has a modest azimuthal brightness asymmetry and a comparatively dim interior. Preferably the dominant model of the accretion disk that reproduces these observations is a magnetically arrested disk with inclination $i\leq 30$ degrees. (Event Horizon Telescope Collaboration, 2022b, a, c, d, e).
@@
The shadow and surrounding photon ring involve a special curve known as the apparent boundary (Bardeen, 1973), and it is well known as the critical curve on the image plane. A light ray with an impact parameter equal to that of the critical curve asymptotically approaches a bound photon orbit when traced backwards from its observation by a distant observer. The existence of such a critical curve produces an observable photon ring: a bright and narrow ring-shaped feature near the critical curve due to the emission of radiation by the optically thin matter from the region the central object, whose photons travel close to the unstable photon orbit before reaching the observer (Jaroszynski & Kurpiewski, 1997; Falcke et al., 1999; Johannsen & Psaltis, 2010).
@@
Additionally, the light rays associated with the photon ring typically orbit the central object multiple times before reaching the observer. This can lead to a significant increase in brightness that may continue to grow without limit in the absence of absorption, as the optical path lengths near the unstable photon orbit become increasingly long; nevertheless, the specific form of the visible photon ring is not influenced by the astronomical source profile (Gralla et al., 2020). Therefore, it can be used as a probe to test general relativity and models for compact objects. In the absence of rotation, one obtains a single unstable circular photon orbit corresponding to an extremum of the effective potential, which is called a photon sphere. When spherical symmetry is lost, the set of unstable photon orbits is deformed and thickened, giving rise to a photon shell (Misner et al., 1973).
@@
Separating the contributions of the background geometry and the influence of the surrounding flow in shadow images has become a critical challenge in the field. In the case of a spherical accretion configuration, the size of the observed shadow for a Schwarzschild black hole is closely influenced only by the spacetime geometry, without being affected by the emission profile (Narayan et al., 2019). However, considering an optically and geometrically thin disk around a compact object, later studies indicate that the black hole image is composed of a dark region and a bright region consisting of several components: direct emission, lensing rings, and photon rings, and the size of the shadow depends on the location of the inner edge of the accretion disk, and the intensity of the observed luminosity (Gralla et al., 2019). Those important features may be regarded as a way to distinguish between black holes and black hole mimickers in the universe (Mazur & Mottola, 2023; Visser & Wiltshire, 2004; Casadio et al., 2024) In fact, the brightness and width of the lensing ring depend on the background geometry and the emission features of the accretion disk surrounding the compact object, and the observed image is dominated by the direct component (Rosa & Rubiera-Garcia, 2022). Nonetheless, this is only true for axial observations, as the direct and lensed contributions lose their circular structure for inclined observations (Rosa et al., 2023; Rosa, 2023; Rosa:2024bqv).
If the spacetime under consideration features an innermost stable circular orbit (ISCO), i.e., a radius below which circular orbits become unstable, this radius accurately predicts where the inner edge of the accretion disk lies (Afshordi & Paczyński, 2003). However, in models of optically thin and geometrically thick accretion disks, for which the inner edge of the disk can penetrate inside the unstable photon orbit, then the minimum size of the shadow is not necessarily limited to the critical curve and can be strongly reduced (Zhang et al., 2024; Vincent et al., 2022).
@@
One of the main advantages of shadow observations is to provide a framework with which to probe deviations with respect to the Kerr or Schwarzschild solutions. However, the limited precision of the first set of observations does not allow for a precise identification of the underlying geometry (Wielgus, 2021; Daas et al., 2023; Lin et al., 2022). Therefore, it is difficult to underestimate theoretical efforts to calculate the forms and properties of shadows produced by black holes or other alternative compact objects in astrophysical environments. Besides these two well-studied Schwarzschild and Kerr hypotheses, at the theoretical level, several alternative models for compact objects have been proposed, featuring a wide range of properties that could potentially fill the gap in the interpretation and understanding of the observational data (Cardoso & Pani, 2019).
In this work, we are interested in studying the shadow properties in the background of a distorted and deformed compact object due to quadrupole moments. This solution is a generalisation of the $\rm q$-metric previously analysed in other publications, e.g. (Erez & Rosen, 1959; Faraji, 2022). We show that moderate changes in the quadrupole moment could lead to significant alterations in the properties of the shadow.
@@
On one hand, we see that the radius of the photon ring, the innermost stable circular orbit (ISCO), and optical appearance can be different depending on the parameters of the geometry and intensity profile. On the other hand, we can deduce information about the source from its observed intensity profile. The latter has a crucial importance in the test of the nature of compact objects and, therefore, in the test of the strong field regime in the general theory of relativity.
@@
Another important issue in astrophysics is the observation of abnormally high velocities in the outer region of galaxies due to the surrounding invisible dark matter (Brand & Blitz, 1993; Sofue & Rubin, 2001). The gravitational tidal forces induced by dark matter influence the propagation of light. Therefore, the presence of dark matter can be analysed through the properties of the shadow via the assumption of different equations of state, e.g. (Hou et al., 2018a, b; Haroon et al., 2019). However, in these studies the results are model-dependent. To explore the effect of dark matter, we consider its behaviour as an effective mass, which leads to modifications in the background geometry, e.g. (Leung et al., 1997; Konoplya & Zhidenko, 2011; Konoplya et al., 2019; Konoplya, 2019).
@@
The paper is organized as follows: in Section 2, we introduce the generalized $\rm q$-metric; in Section 3 we review the geodesic motion of timelike and null particles in this background, we introduce the accretion disk models, and we produce the observed shadow images and intensity profiles; in Section 4 we introduce the dark matter model as a modification of the radial mass function and obtain the corresponding observed shadow images and intensity profiles; and in Section 5 we summarize our results and trace prospects for future work. Throughout the paper, we use a system of geometric units in which $G=1=c$ unless stated otherwise.
2 Generalized $\rm q$-metric spacetime
@@
One of the generalisations of the Schwarzschild spacetime recurring to the Weyl’s family of solutions (Weyl, 1917) is the $\rm q$-metric. This asymptotically flat solution of the Einstein field equations represents the exterior gravitational field of an isolated static and axisymmetric mass distribution containing a quadrupole moment, and it can be used to investigate the exterior fields of slightly deformed astrophysical objects in the strong-field regime (Quevedo, 2011). This metric has been further generalised by relaxing the assumption of an isolated compact object, or equivalently asymptotic flatness, by considering the presence of an exterior distribution of mass in its vicinity (Faraji, 2022). Since this spacetime is not asymptotically flat by construction, the solution presented here is a vacuum solution of the Einstein field equations in the region near the central deformed object. The extension of this vacuum region depends on the chosen values of the parameters—smaller parameter values lead to a more extended vacuum region. Beyond this vacuum region, the spacetime could be extended by introducing a proper stress-energy tensor, which can be interpreted as a general external distribution of matter. Such a generalisation resembles a magnetic surrounding (Ernst, 1976) and it is also characterised by multipole moments. However, due to the subdominant effects of the higher multipole moments, it is appropriate to consider only the contributions to the metric up to the quadrupole moment for analyses in an astrophysical context. Therefore, in this spacetime, there are two independent quadrupole moments: one responsible for the deformation of the central object, $\alpha$, and the other associated with the relaxation of asymptotic flatness, linked to the concept of an external matter distribution, $\beta$. Here we adopt a set of coordinates analogous to the usual spherical coordinates $\left(t,r,\theta,\phi\right)$ in spherically symmetric spacetimes [^1] In this spacetime, the coordinate $r$ represents a radial parameter where $r=\text{const.}$. hypersurfaces are topologically spheres. However, due to the deformation of spacetime, the geometric properties of these hypersurfaces, such as surface area and curvature, are altered. This modification is reflected in the metric components, which adjust the standard interpretation of distances and angles in spherical coordinates.. the line element describing this generalized $\rm q$-metric is written as
@@
where $M$ is a parameter that plays the role of the mass of the central object, and the metric functions $\hat{\psi}$ and $\hat{\gamma}$ that control the distortion of the spacetime also up to quadrupole read as follows
@@
where $\alpha$ and $\beta$ are quadrupole parameters. The coordinates range in the intervals $t\in(-\infty,+\infty)$, $r\in(0,+\infty)$, $\theta\in[0,\pi]$, and $\phi\in[0,2\pi)$. Under the assumptions outlined above, this metric contains three free parameters: the deformation parameter $\alpha$, the distortion parameter $\beta$, and the total mass $M$ that can be absorbed through a redefinition of the radial coordinate. In the equatorial plane, described by $\theta=\pi/2$, the metric functions reduce to
@@
$\hat{\gamma}$ $=\beta\left[-2(1+\alpha)\left(\frac{r}{M}-1\right)+\frac{\beta r}{4M}\left(\frac{r}{M}-2\right)\left(\left(\frac{r}{M}-1\right)^{2}-1\right)\right]\,.$
In fact, the presence of a quadrupole can have a notable impact on the geometric properties of space-time when compared to the Schwarzschild solution, see e.g. (Faraji & Trova, 2021, 2022). Mathematically, the deformation parameter is bounded by $\alpha\in(-1,\infty)$. However, in an appropriate astrophysical context, other studies see e.g. (Faraji & Trova, 2021, 2022), have demonstrated that the most relevant domain for $\alpha$ is almost $(-1,1)$. It is important to note that geodesics are influenced by both parameters $\alpha$ and $\beta$ simultaneously. In fact, calculating the radius of the innermost stable circular orbit (ISCO) and of the unstable photon orbit, which is done by setting the second and first derivatives of the respective effective potentials to zero, leads to a relationship between $\alpha$ and $\beta$. This happens because, for a given value of $\alpha$ the range of the parameter $\beta$, specifically in the equatorial plane, depends on $\alpha$ and arises directly from solving the mentioned equations. In addition, the range of the parameter $\beta$ turns out to be significantly smaller than the latter. Additionally, in this work, we are interested in configurations that are approximately spherically symmetric, and consequently, the quadrupole parameters should be relatively small in magnitude. To guarantee that this approximation of spherical symmetry is valid, one needs to keep the absolute values of the parameters $\alpha$ and $\beta$ below the values $0.01$ and $10^{-6}$, respectively, which guarantees that the deviations from spherical symmetry of the $g_{tt}$, $g_{rr}$, and $g_{\phi\phi}$ components of the spacetime metric are always below $1\%$ in the region where photons propagate (see Appendix A). This solution is the exact superposition of an oblate or prolate compact object, surrounded by an external distribution of matter. Recently, the impact of superposed fields in the motion of free test particles in this spacetime was studied (Faraji, 2022), and it was shown how the positions of important astrophysical orbits, such as ISCO for particles, and photon circular orbits, depend on the parameters of the metric. In Figure 1, we plot the second derivative of the particle potential $V_{\rm part}=-g_{tt}\left(\frac{L^{2}}{r^{2}}+1\right)$, where $L$ is the angular momentum of the particle, and the first derivative of the photon effective potential $V_{\rm phot}=-\frac{g_{tt}}{r^{2}}$. The results indicate that an increase in $\alpha$ shifts the inflection points of the particle effective potential and the extrema of the photon effective potential to larger radial coordinates. In contrast, an increase in $\beta$ causes the behaviour of these two features to shift in opposite directions due to the effect of external distribution. In this case, the unstable photon orbit, being closer to the object than ISCO, feels the gravitational pull of the central object more strongly than the influence of the external fields. For instance, if $\beta>0$ the external quadrupole tends to push the unstable photon orbit inward, increasing the curvature near the central object. The ISCO, located farther out, feels a weaker pull from the central object but a stronger influence from the external quadrupole, which causes it to shift outward.
Figure 1: Second radial derivative of the particle effective potential $V^{\prime\prime}_{\rm part}$ (top row), and first radial derivative of the photon potential $V^{\prime}_{\rm phot}$, both as a function of the radius $r$, with $\beta=0$ and varying $\alpha$ (left panels) and with $\alpha=0$ and varying $\beta$ (right panels), at the equatorial plane $\theta=\frac{\pi}{2}$. The radius of the ISCO is defined as the radius at which $V^{\prime\prime}_{\rm part}=0$, and the radius of the photon orbit is defined as the radius at which $V^{\prime}_{\rm phot}=0$. We observe that a change in $\alpha$ causes the radii of the ISCO and the photon orbit to change in the same direction, whereas a change in $\beta$ causes these radii to change in opposite directions.
Moreover, from an astrophysical perspective, the negative values of the quadrupole parameters are rather counter-intuitive since they describe prolate compact objects, while positive values correspond to oblate compact objects that are widely observed in the universe.
3 Shadows and light rings